SINGLE QUANTUM-STATE RESOLVED COLLISION RELAXATION CROSS SECTION OF HIGHLY VI-BRATIONALLY EXCITED SO_{2} MEASURED BY KINETIC QUANTUM BEAT SPECTROSCOPY

\underline{MIN ZHANG}, JIANQIANG MA, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104; HAI-LUNG DAI, Department of Chemistry, Temple University, Philadelphia, PA 19122.

The collision relaxation cross section of highly vibrationally excited SO_{2} at energies around 45,000 cm^{1}, which is just below the dissociation limit, has been measured for several single, highly vibrationally excited levels. The measurement was conducted on rotationally cooled SO_{2} in supersonic jet condition using quantum beat spectroscopy. Quantum beat in laser induced fluorescence from the C̃ state is due to vibronic coupling between zero-order rovibronic levels of C̃ and X̃ states and can be analyzed to give the decay rate constants of the high vibrational levels of the X̃ state. It is found that the collision relaxation cross section of the single highly vibrationally excited level shows a strong correlation with the corresponding vibronic coupling matrix element, likely reflecting that the part of the vibrational character of the single high vibrational level that gives a larger coupling with the C̃ state also provide a stronger transition dipole moment contribution to its collision relaxation.